### v1

# Empirical Analysis of the Model-Free Valuation Approach: Hedging Gaps, Conservatism, and Trading Opportunities

Abstract.
In this paper we study the quality of model-free valuation approaches for financial derivatives by systematically evaluating the difference between model-free super-hedging strategies and the realized payoff of financial derivatives using historical option prices from several constituents of the S&\&$500$ between $2018$ and $2022$.
Our study allows in particular to describe the realized gap between payoff and model-free hedging strategy empirically so that we can quantify to which degree model-free approaches are overly conservative. Our results imply that the model-free hedging approach is only marginally more conservative than industry-standard models such as the Heston-model while being model-free at the same time.
This finding, its statistical description and the model-independence of the hedging approach enable us to construct an explicit trading strategy which, as we demonstrate, can be profitably applied in financial markets, and additionally possesses the desirable feature with an explicit control of its downside risk due to its model-free construction preventing losses pathwise.
Keywords: Model-free price bounds, Hedging, Linear programming, Profitable Trading
August 9, 2025
1National University of Singapore, Department of Mathematics,
21 Lower Kent Ridge Road, 119077.
1. Introduction
Building on the foundational works beiglbock2013model; beiglbock2016problem; henry2016explicit, which establish the connection between solutions of the martingale optimal transport problem, model-free price bounds for financial derivatives and associated hedging strategies, the literature on model-free approaches and their application to derivative valuation has grown rapidly and still continues to grow. See, for instance, ansari2024improved; cheridito2017duality; chiu2023model; davis2016model; dolinsky2014martingale; de2020bounds; fahim2016model; hobson2012robust; hobson2015robust; eckstein2021robust; lux2017improved, among many others. Using only available market data of liquidly traded put- and call options, the aforementioned model-free approaches compute price bounds for less liquidly traded (so called exotic) options solely under the paradigm of no arbitrage and of consistency with observed liquid option prices but, importantly, without using any assumptions on the stochastic dynamics of the underlying assets. The resultant price bounds are then calculated as extreme prices among all calibrated arbitrage-free models, and therefore can be understood to be model-free and to respect the so called Knightian uncertainty (knight1921risk), i.e., the unquantifiable risk of having chosen a wrong model.
While appealing from a risk management perspective, not imposing model assumptions theoretically comes at the application-prohibitive cost of resultant model-free price bounds by construction being larger than the price in any calibrated model that is also consistent with liquid option prices and arbitrage-free, making the approach less appealing from a practitioner’s perspective, see also the discussions in eckstein2021martingale; lutkebohmert2019tightening; dolinsky2018super.
To the best of our knowledge there exists however little to no research providing empirical evidence for the actual conservativeness of model-free approaches, i.e., research answering the question how much the corresponding prices bounds are too wide in real financial markets, and whether they indeed prevent applicability. According to the duality theory developed in beiglbock2013model; beiglbock2016problem; henry2016explicit, to study the quality of model-free upper price bounds, one can equivalently study prices of model-free super-replication strategies (or sub-replication strategies for lower price bounds). Namely, if the gap between the realized payoff of a price bound-attaining model-free super-replication strategy and the realized derivative payoff is (too) large, this indicates that the strategy’s price and hence its initial cash endowment was too large as well and could be reduced to obtain a less conservative strategy with payoff closer to the derivative’s actual payoff.
Relying on the above outlined idea, the present paper aims at closing an existing research gap by studying systematically the difference between model-free hedging strategies (corresponding to model-free price bounds) and realized payoffs of exotic derivatives. For our empirical study, we use historical option prices written on seven of the largest assets by market capitalization of the S&\&$500$ between $2018$ and $2022$. This paper is the first to empirically quantify the realized hedging performance and conservatism of model-free super-hedging strategies using real market data. Our findings challenge the prevailing common belief that model-free methods are impractically conservative, at least if the involved assets and sufficiently many call- and put options are liquidly traded in the market, as it is the case for our study.
Our analysis shows that model-free hedging strategies lead to a hedging performance lacking not much behind the ones of industry-standard models such as the Black-Scholes-, Heston- and Dupire’s local volatility model. As, in contrast to the aforementioned industry-standard models, model-free super hedging strategies do not rely on a specific probabilistic model of the underlying security, they always lead to larger payoffs than the derivative given any realization of the security. We use this robustness property to propose a trading strategy that makes use of both our statistical description of the payoff gap and the model-free construction of the hedging strategy. Trading for this strategy is triggered whenever observed market quotes for the derivative are close enough to its model-free price bounds where the degree of proximity determines the probability of a positive payoff which can be exactly described according to our statistical analysis. The strategy is robust in the sense that up to the initial investment no additional money can be lost due to its model-free construction, consequently leading to a low-risk trading strategy with an easy-to-understand risk profile.
The remainder of this paper is as follows. In Section 2 we provide an overview on the used data, Section 3 describes in detail the employed mathematical approach to compute the super-hedging strategies based on linear programming. Section 4 presents the results of our analysis when being applied to data, and Section 5 discusses and analyzes a trading strategy that is based on our statistical findings.
2. Data and technology
2.1. Data
For our study, we use data received from Wharton Research Data Services. The data comprises bid and ask prices of traded put- and call options in the time between Jan 1 2018 and Dec 31 2022. The considered underlying securities on which the options are written are the stocks of Amazon.com Inc, JP Morgan Chase & Co., Johnson & Johnson, Procter & Gamble Co., Walmart Inc, VISA Inc, and Tesla Inc.
2.2. Code
For the sake of transparency and replicability, we provide the used Python-code in an accompanying github-repository [^1] The repository can be found under https://github.com/QIYIHAN/Empirical-Price-Bounds.. Unfortunately, legal constraints prevent us from providing the historical option price data in the github repository.
3. Methodology
Our methodology consists mainly of three steps. In the first step we pre-process the data, second we compute model-free price bounds and corresponding hedging strategies via linear programming, and third we compute the gap between the realized payoff of an exotic financial derivative and the payoff of the previously computed hedging strategy given outcomes of the underlying assets described in Section 2.1. In the following we introduce the mathematical setup in Section 3.1 and detail all the computational steps in Section 3.2 and 3.3.
3.1. Model-free valuation of financial derivatives
We consider some financial derivative providing a payoff $\Phi(S_{t_{1}},\dots,S_{t_{n}})$ at some future time $t_{n}>0$ with $S_{t}$ denoting the value of the underlying asset at time $t$. To determine a range of model-free prices for $\Phi$, we compute the minimal super-replication price of $\Phi$ when investing in so called semi-static trading strategies (see, e.g. davis2007range for more details on semi-static trading), i.e., trading strategies investing once in liquidly traded European call options and adjusting dynamically its investment in the underlying security. These strategies possess a payoff at time $t_{n}$ of the form
(3.1) $u_{d,(\theta_{i,j}),(H_{i})}:\mathbb{R}^{n}$ $\rightarrow\mathbb{R}$
$(S_{t_{1}},\dots,S_{t_{n}})$ $\mapsto e^{r_{(t_{n}-t_{0})}\cdot(t_{n}-t_{0})}d$
$+\sum_{i=1}^{n}\sum_{j=1}^{M_{i}}e^{r_{(t_{n}-t_{i})}\cdot(t_{n}-t_{i})}(\theta_{i,j}^{\rm call,+}-\theta_{i,j}^{\rm call,-})(S_{t_{i}}-K_{i,j}^{\rm call})^{+}$
$+\sum_{i=0}^{n-1}H_{i}(S_{t_{1}},\dots,S_{t_{i}})(S_{t_{i+1}}-S_{t_{i}}e^{r_{(t_{i+1}-t_{i})}\cdot(t_{i+1}-t_{i})}),$
where $r_{\tau}\in\mathbb{R}$ denotes the interest rate used for discounting of a time period of length $\tau$, $d\in\mathbb{R}$ denotes the cash position, $(\theta_{i,j}^{\rm call,+},\theta_{i,j}^{\rm call,-})\in\mathbb{R}_{\geq 0}^{2}$ the holdings of the call options, $K_{i,j}^{\rm call}\in\mathbb{R}_{\geq 0}$ the strikes of the call options, $M_{i}\in\mathbb{N}$ the number [^2] Note that we do not need to distinguish between the number of strikes considered for the call options since we can simply use the larger of the two and include options multiple times if necessary to match the number of options considered. of the call options with maturity $t_{i}$, $H_{i}(S_{t_{1}},\dots,S_{t_{i}})\in\mathbb{R}$ (while $H_{0}$ is a constant) denote the amount of the investment in the underlying security at time $t_{i}$.
Remark 3.1.
  • (i)
    As our analysis involves infrequent rebalancing of the dynamic positions $(H_{i})_{i}$, the effects of transaction costs for $(H_{i})_{i}$ are expected to be rather small, and thus, even though mathematically feasible, we have decided to not include assumptions on transaction costs related to trading in the underlying asset $S$ in our trading strategy. How to mathematically model semi-static trading that includes transaction costs is $S$ is described, e.g., in (cheridito2017duality, Section 3.1).
  • (ii)
    For the sake of presentation we have decided to not include put options into (3.1) which is well justified by the put-call parity (stoll1969relationship) and since the call options we consider are sufficiently liquid by our data selection process described in Section 3.3.1.
To compute the price of the strategy with payoff $u_{d,(\theta_{i,j}),(H_{i})}$, we denote the ask and bid prices of the considered call options by $h_{i,j}^{\rm call,+}$ and $h_{i,j}^{\rm call,-}$ respectively. Then, to compute the largest model-free price for $\Phi$ which does not allow for model-free arbitrage (in line with the arbitrage notion from (acciaio2016model, Definition 1.2)), we compute the super-hedging price by
(3.2) $\overline{\operatorname{P}}(\Phi)=\inf_{d,(\theta_{i,j}),(H_{i})}\Bigg{\{}$ $d+\sum_{i=1}^{n}\sum_{j=1}^{M_{i}}(\theta_{i,j}^{\rm call,+}h_{i,j}^{\rm call,+}-\theta_{i,j}^{\rm call,-}h_{i,j}^{\rm call,-})\leavevmode\nobreak\$
$\text{ subject to}\leavevmode\nobreak\ u_{d,(\theta_{i,j}),(H_{i})}(S)\geq\Phi(S)\text{ for all }S\in\mathbb{R}_{\geq 0}^{n}\Bigg{\}}$
compare also the expositions in the related literature, e.g., in acciaio2016model, (neufeld2021model, Section A.1) or (neufeld2022deep, Section II). Correspondingly, a sub-hedging price can be computed via
(3.3) $\underline{\operatorname{P}}(\Phi):=-\overline{\operatorname{P}}(-\Phi).$
By construction, market prices outside the interval $\underline{\operatorname{P}}(\Phi),\overline{\operatorname{P}}(\Phi)]$ allow to pocket in a model-free arbitrage profit through investing in the corresponding semi-static sub- or super-hedging strategies with price attaining the bounds of $\underline{\operatorname{P}}(\Phi),\overline{\operatorname{P}}(\Phi)] -$ independent of the realized outcome of the underlying security.
Remark 3.2.
Note that if a financial model is calibrated to observed bid-ask prices of call options, then arbitrage-free prices of an exotic derivative $\Phi$ can be computed by $\mathbb{E}_{\mathbb{Q}}\left[e^{-r_{t_{n}-t_{0}}(t_{n}-t_{0})}\Phi(S_{t_{1}},\dots,S_{t_{n}})\right]$ where $\mathbb{Q}$ is a martingale measure that fulfils $\mathbb{E}_{\mathbb{Q}}\left[e^{-r_{t_{i}-t_{i}}(t_{i}-t_{0})}(S_{t_{i}}-K_{i,j}^{\rm call})^{+}\right]\in\left[h_{i,j}^{\rm{call},-},\leavevmode\nobreak\ h_{i,j}^{\rm{call},+}\right]$ for all $i,j$. If we denote the set of all such probability measures by $\mathcal{Q}$, then (cheridito2017duality, Section 3.1.3) or (neufeld2021model, Section A.1) show that under mild assumptions
$\overline{\operatorname{P}}(\Phi)=\sup_{\mathbb{Q}\in\mathcal{Q}}\mathbb{E}_{\mathbb{Q}}\left[e^{-r_{t_{n}-t_{0}}(t_{n}-t_{0})}\Phi(S_{t_{1}},\dots,S_{t_{n}})\right],$
i.e., the smallest super-replication price $\overline{\operatorname{P}}(\Phi)$ can also be understood as the largest price among all arbitrage-free financial models calibrated to market data.
3.2. Computation of model-free super-hedges with linear programming
For the computation of model-free super-hedging price $\overline{\operatorname{P}}(\Phi)$ as defined in (3.2) and the associated super-hedging strategy we mainly rely on a numerical routine using linear programming (LP) as outlined in similar settings in (henry2013automated, Section 3) and guo2019computational. Since the computation of the sub-hedging strategy follows analogously through using the relation in (3.3), we focus with our presentation on the computation of a super-hedging strategy at minimal price $\overline{\operatorname{P}}(\Phi)$ in the two time-step case $n=2$, i.e.,
(3.4) $\overline{\operatorname{P}}(\Phi)=\inf_{d,(\theta_{i,j}),H_{0},H_{1}}\Bigg{\{}$ $d+\sum_{i=1}^{2}\sum_{j=1}^{M_{i}}(\theta_{i,j}^{\rm call,+}h_{i,j}^{\rm call,+}-\theta_{i,j}^{\rm call,-}h_{i,j}^{\rm call,-})$
$\text{ subject to}\leavevmode\nobreak\ u_{d,(\theta_{i,j}),(H_{0},H_{1})}(S)\geq\Phi(S)\text{ for all }S\in\mathbb{R}_{\geq 0}^{2}\Bigg{\}}.$
The above problem (3.4) falls in the class of infinite-dimensional linear programming problems. To solve (3.4) numerically, we aim at transforming (3.4) to a finite dimensional problem. To this end, we first consider a (sufficiently large) grid
$\mathbb{G}=(\mathbb{G}^{1},\mathbb{G}^{2})\subset\mathbb{R}_{\geq 0}^{2},$
where both $\mathbb{G}^{1}$ and $\mathbb{G}^{2}$ are one-dimensional grids with $N$ meshes that discretize the values attained by the distributions of $S_{t_{1}}$ and $S_{t_{2}}$, respectively. Given a function $H_{1}:\mathbb{R}\rightarrow\mathbb{R}$, representing the dynamic investing in the underlying security, as defined in (3.1), for each grid point $S_{t_{1}}^{\ell}\in\mathbb{G}^{1}$, we define the value $H_{1}^{\ell}:=H_{1}(S_{t_{1}}^{\ell})$. Finally, to approximate $H_{1}$ on the whole reals, we apply a cubic spline interpolation (see, e.g., ahlberg2016theory) using $(H_{1}^{\ell})_{\ell=1,\dots,N}$, and obtain a cubic spline function
$\rm{spline}\left((H_{1}^{\ell})_{\ell=1,\dots,N}\right):\mathbb{R}\rightarrow\mathbb{R}.$
which coincides with $H_{1}$ on $\mathbb{G}^{1}$ and interpolates smoothly the non-grid values. The above outlined procedure allows us to define an approximating finite dimensional linear programming problem [^3] Note that for the formulation in (3.5), we do not necessarily need a smooth function approximation as only the grid points on $\mathbb{G}$ are considered for the super-replication condition. However, in later sections we will use the corresponding cubic-spline function to represent the learned strategy also applied to values off the grid.
(3.5) $\widetilde{\overline{\operatorname{P}}}(\Phi,\mathbb{G})=\inf_{d,(\theta_{i,j}),H_{0},(H_{1}^{\ell})_{\ell}}\Bigg{\{}$ $d+\sum_{i=1}^{2}\sum_{j=1}^{M_{i}}(\theta_{i,j}^{\rm call,+}h_{i,j}^{\rm call,+}-\theta_{i,j}^{\rm call,-}h_{i,j}^{\rm call,-})$
$\text{ subject to }\leavevmode\nobreak\ u_{d,(\theta_{i,j}),(H_{0},\rm{spline}\left((H_{1}^{\ell})_{\ell}\right)}(S)\geq\Phi(S)\text{ for all }S\in\mathbb{G}\Bigg{\}}.$
To find numerical solutions to $\widetilde{\overline{\operatorname{P}}}(\Phi,\mathbb{G})$, we apply a cutting plane algorithm as outlined in (henry2013automated, Section 3.1) of which we provide details in the appendix, compare Algorithm 1.
3.3. Computation of the gap between the realized payoff and the payoff of the hedging strategy
This section outlines the method employed to compute the difference between the outcomes of minimal model-free super-replication strategies (as well as maximal model-free sub-replication strategies) and the underlying derivative payoff.
3.3.1. Preprocessing of the data
From the option price data described in Section 2.1 we first select all pairs of option prices with the same underlying stock and fixed length of maturity difference $t_{2}-t_{1}>0$, where $t_{1},t_{2}$ are two futures times relevant for the option payoff. To obtain sufficiently reliable prices, for each fixed maturity difference, we select the $20$ most liquid call options (and associated strikes) by choosing those with the highest trading volume. Then, to avoid arbitrage among the considered options (in particular butterfly arbitrage, see gatheral2014arbitrage), we apply the arbitrage repair method introduced in Cohen2020 to the mid-prices. Using the repaired mid price $C^{*}$ and original spread $\delta=C_{\rm ask}-C_{\rm bid}$, we obtain eventually repaired bid and ask prices $C_{\rm bid}^{*}$ and $C_{\rm ask}^{*}$ via the relations
$C_{\rm ask}^{*}=C^{*}+\frac{1}{2}\delta,\qquad C_{\rm bid}^{*}=C^{*}-\frac{1}{2}\delta.$
Further details of the preprocessing procedure are described in the accompanying github repository which can be found under https://github.com/QIYIHAN/Empirical-Price-Bounds.
3.3.2. Computing interest rates
To model the interest rate $(r_{t})_{t}$ we use throughout our study a Nelson–Siegel model (Nelson-Seigel) fitted to US treasury yields with the historical parameters that can be found under https://www.federalreserve.gov/data/nominal-yield-curve.htm.
3.3.3. Computation of the gap
Upon computing $\widetilde{\overline{\operatorname{P}}}(\Phi,\mathbb{G})$ defined in (3.5), by using Algorithm 1, we obtain the optimal values of the parameters $d$, $(\theta_{i,j})_{i=1,2,\atop j=1,\dots,M_{1}}$, $H_{0}$ and $(H_{1}^{\ell})_{\ell=1,\dots,N}$, and interpolate $(H_{1}^{\ell})_{\ell=1,\dots,N}$ by a cubic spline interpolator $\widetilde{H}_{1}:=\rm{spline}\left((H_{1}^{\ell})_{\ell=1,\dots,N}\right)$. Denoting the realized prices of the underlying security at times $t_{1}$ and $t_{2}$, respectively by $S\in\mathbb{R}^{2}$,we compute the raw value of the gap between realized hedging strategy and realized derivative payoff by
$u_{d,(\theta_{i,j}),(H_{0},\widetilde{H}_{1})}(S)-\Phi(S).$
Eventually, to make the gap values comparable across different underlying securities, we normalize the raw gap by the price of the underlying security at the initial time $t_{0}$ denoted by $S_{0}$, i.e., we define the normalized hedging gap (or simply gap in the following) by
(3.6) $\operatorname{gap}=\frac{u_{d,(\theta_{i,j}),(H_{0},\widetilde{H}_{1})}(S)-\Phi(S)}{S_{0}}.$
4. Empirical Tightness of Model-Free Hedging
In this section we provide and discuss the results of the presented numerical routine from Section 3 applied to the data presented in Section 2.
4.1. The payoff gap among different time-to-maturities
We start our analysis by considering a forward start at-the-money call option possessing a payoff of the form $\Phi(S_{t_{1}},S_{t_{2}}):=(S_{t_{2}}-S_{t_{1}})^{+}:=\max\{S_{t_{2}}-S_{t_{1}},0\}$, where $S_{t_{1}}$ and $S_{t_{2}}$ denote the value of the underlying asset at future times $t_{1}$ and $t_{2}$, respectively, i.e., this is a call option where the strike is determined at time $t_{1}$ such that the option is at-the-money. Using the data described in Section 2.1, we evaluate the gaps, computed according to (3.6), between the realized payoff of the minimal model-free super-hedging strategy and that of the forward start option across different time-to-maturities, i.e., for different distances between $t_{0}$ (the time when the option contract is set up) and $t_{1}$, while the difference $t_{2}-t_{1}$ is fixed to $28$ days. We illustrate median and different quantile ranges of the computed gaps as well as the underlying sample size in Figure 1. We observe that with the initial date $t_{0}$ approaching the date $t_{1}$, i.e., the distance between $t_{0}$ and $t_{1}$ decreasing, the gaps tend to become smaller: the median of the payoff gap is decreasing, and the quantile ranges are shifting towards smaller gap values. Similarly, the variance of the gap becomes smaller for a decreasing time-to-maturity corresponding to the empirical fact that the farther the maturity lies in the future the higher the associated uncertainty of the outcome of the asset and the payoff (measured in volatility). These observations become more unstable with an increasing time-to-maturity, which we attribute mainly to a smaller available sample size for longer maturities, indicated by the dashed orange line in Figure 1.
Figure 1. The different quantile ranges of the payoff gap for super-hedging with payoff function $(S_{t_{2}}-S_{t_{1}})^{+}$.
4.1.1. Sub-hedging
In addition to the case of super-hedging discussed above, we also study the gap between realized outcomes of model-free sub-hedging strategies and the realized payoff of a forward start option with payoff $(S_{t_{2}}-S_{t_{1}})^{+}$. The statistics (median, quantile ranges) of the gap values under sub-hedging are illustrated in Figure 2. We remark that according to (3.6) the gap values in this situation are by construction mostly negative, but, in contrast to intuition, there are some mildly positive gap values in Figure 2 which we explain through the absence of transaction costs and due to the discretization inherent in Algorithm 1 in which we accept the result as long as the gap values do not exceed the tolerance level $\rm TOL$}.
The observations in Figure 2 are mostly in line with those from Figure 1, although the median of the gap appears to be more stable. Overall, as in Figure 1 the empirical payoff gap widens with an increasing time-to-maturity, where the gap for the model-free sub-hedging strategy widens less significantly than the model-free super-hedging strategy in this case.
Figure 2. The different quantile ranges of payoff gap for sub-hedging with payoff function $(S_{t_{2}}-S_{t_{1}})^{+}$.
4.2. The payoff gap among different payoffs
We repeat the experiments from Section 4.1 with several other payoff functions. The corresponding statistics, shown in Figures 35, are consistent with the findings reported in Section 4.1.
Figure 3. The different quantile ranges of the payoff gap for super-hedging with payoff function $\frac{S_{t_{1}}+S_{t_{2}}}{2}-\sqrt{S_{t_{1}}\cdot S_{t_{2}}}$.
Figure 4. The different quantile ranges of the payoff gap for super-hedging with payoff function $\log(S_{t_{2}}/S_{t_{1}})$.
Figure 5. The different quantile ranges of the payoff gap for super-hedging with payoff function $\sqrt{S_{t_{1}}\cdot S_{t_{2}}}$.
Moreover, in the appendix, in Figures 11 - 14, we illustrate the distribution of the gap values for all payoffs under both super-hedging and sub-hedging by depicting the respective histograms.
4.3. Comparison with model-based hedges
Next, to set the size of the payoff gaps reported in Section 4.1 and Section 4.2 in relation, we compare the quality of the computed model-free super- and sub-hedging payoffs with hedging payoffs computed under several industry-standard models that are calibrated to the option data presented in Section 2.1. To keep the paper self-contained, we present the considered benchmark models briefly in Section D. The corresponding hedging strategies in each of the models are implemented by considering a delta-hedging strategy with adjustment to the respective Gamma and Theta sensitivities, compare for more details, e.g., hull2017optimal. As the results from Section 4.1 have revealed, the gap distribution of model-free hedges under different derivatives turns out to be similar. Hence, for the analysis in this section we focus solely on the forward-start call option with payoff $(S_{t_{2}}-S_{t_{1}})^{+}$. From initiation date $t_{0}$ to $t_{2}$, we compute the option delta with on a daily basis, compare Section D for more details.
Figure 6. The distribution of the gap between derivative payoff and different hedging strategies for a derivative payoff of the form $(S_{t_{2}}-S_{t_{1}})^{+}$ in model-based and model-free hedging approaches.
Table 1. Summary Statistics for different hedging strategies for a derivative payoff of the form $(S_{t_{2}}-S_{t_{1}})^{+}$ using either model-based or model-free hedging approaches.
Black-Scholes Local Volatility Heston Super-Hedging Sub-Hedging
Mean 0.00 -0.01 0.02 0.05 -0.03
Std.dev. 0.10 0.09 1.37 0.93 0.07
Min -1.47 -1.47 -1.32 -0.04 -1.42
Max 0.38 0.36 171.98 217.95 0.06
25 % -0.02 -0.03 -0.02 0.01 -0.04
75 % 0.05 0.03 0.05 0.06 -0.00
In Figure  6 and Table 1, we compare the distribution of the resultant hedging errors (defined as in (3.6)) of the model-based hedging strategies with the hedging errors of the model-free super and sub-hedging strategies among all derivatives whose payoffs are $(S_{t_{2}}-S_{t_{1}})^{+}$ where the difference between $t_{1}$ and $t_{2}$ is fixed to $28$ days, and where the hedges are evaluated on the data described in Section 2.1. We observe that the hedges under all considered models lead to reasonably tight hedging gaps with interquartile ranges of [0.02,0.05],[0.03,0.03][-0.02,0.05],[-0.03,0.03], and [0.02,0.05][-0.02,0.05], respectively. The main observation is though that the interquartile ranges of the gaps of model-free super-hedges and sub-hedges are given by $[0.01,0.06]$, and $[0.00,0.04]$, respectively and therefore they are of similar size as the model-based hedges, contradicting the intuition and common belief that model-free approaches lead to overly conservative hedging outcomes. As by pricing-hedging duality (see, e.g., eberlein2008duality, hou2018robust), prices of hedges correspond to arbitrage-free prices of the corresponding derivative, the prices produced by model-free pricing approaches turn out to be competitive and rather close to prices produced by industry standard models (and thus traded prices of OTC derivatives) - an observation that we will use in the next section to explicitly construct a profitable trading strategy.
5. A trading strategy based on model-free price bounds
The analysis from Section 4 allows to derive a trading strategy with an easy-to-understand risk control regarding its downside risk. The idea behind the strategy is to exploit discrepancies between market prices and model-free super- or subhedging price bounds: the study from Section 4.3 shows that price bounds of model-free approaches are relatively close to model-based prices of industry standard models; if now indeed the observed price of a derivative lies sufficiently close to these bounds, a model-free hedge is constructed whose payoff dominates the derivative, thereby allowing the investor to lock in a non-negative payoff exceeding the limited initial investment with a high probability. By tuning a threshold parameter (that we later call $c$), the investor can control the trade-off between risk and expected return.
5.1. Description of the strategy
To construct the strategy, let $\overline{u_{d,(\theta_{i,j}),(H_{i})}}$ denote a hedging strategy of the form described in (3.1) attaining the upper price bound $\overline{\operatorname{P}}(\Phi)$ [^4] We can always find such a strategy up to an arbitrarily small tolerance $\varepsilon>0$ which we ignore here for our exposition.. Then, we have by construction of the model-free super-hedge that $\overline{u_{d,(\theta_{i,j}),(H_{i})}}(S)-\Phi(S)\geq 0$ for all $S\in\mathbb{R}^{n}$. Moreover, for all levels of relative hedging gaps $c\in\mathbb{R}$ there exists an associated probability $\overline{q}(c)\in[0,1]$ such that the empirical probability of a realized relative gap larger than $c$ is $\overline{q}(c)$, i.e.,
$\widehat{\mathbb{P}}\left(\frac{\overline{u_{d,(\theta_{i,j}),(H_{i})}}(S)-\Phi(S)}{S_{0}}\geq c\right)=\overline{q}(c),$
where $\widehat{\mathbb{P}}$ denotes the empirical distribution derived from historical stock data $(S^{(i)})_{i=1,\dots,N}=(S_{t_{1}}^{(i)},S_{t_{2}}^{(i)})_{i=1,\dots,N}$ of the respective stock, i.e.,
$\widehat{\mathbb{P}}=\frac{1}{N}\sum_{i=1}^{N}\delta_{(S_{t_{1}}^{(i)},S_{t_{2}}^{(i)})},$
for $\delta_{x}$ denoting the Dirac measure at point $x\in\mathbb{R}^{2}$ and where $N$ denotes the number of historical stock prices used for the construction of $\widehat{\mathbb{P}}$. In the same manner, we denote by $\underline{u_{d,(\theta_{i,j}),(H_{i})}}$ a strategy attaining the lower price bound $\underline{\operatorname{P}}(\Phi)$. Then, for all $c\in\mathbb{R}$ there exists some $\underline{q}(c)\in[0,1]$ such that
$\widehat{\mathbb{P}}\left(\frac{\Phi(S)-\underline{u_{d,(\theta_{i,j}),(H_{i})}}(S)}{S_{0}}\geq c\right)=\underline{q}(c).$
If now an observed market price $\mathfrak{M}(\Phi)$ for the derivative $\Phi$ is close to either $\overline{\operatorname{P}}(\Phi)$ or $\underline{\operatorname{P}}(\Phi)$, i.e., given some $c>0$, we have
(5.1) $\min\{\mathfrak{M}(\Phi)-\underline{\operatorname{P}}(\Phi),\leavevmode\nobreak\ \leavevmode\nobreak\ \overline{\operatorname{P}}(\Phi)-\mathfrak{M}(\Phi)\}\leq c\cdot S_{0},$
then we may invest in $\overline{u_{d,(\theta_{i,j}),(H_{i})}}(S)-\Phi(S)$, or $\Phi(S)-\underline{u_{d,(\theta_{i,j}),(H_{i})}}(S)$ (we choose whatever is the binding condition in (5.1)). For such an investment we pay less than $c\cdot S_{0}$ to subsequently hold a strategy which leads under $\widehat{\mathbb{P}}$ with a probability of either $\overline{q}(\tfrac{\overline{\rm{P}}(\Phi)-{\mathfrak{M}}(\Phi)}{S_{0}})\geq\overline{q}(c)$ or $\underline{q}(\tfrac{{\mathfrak{M}}(\Phi)-\underline{\rm{P}}(\Phi)}{S_{0}})\geq\underline{q}(c)$ to a positive profit. Moreover, and this makes the investment extremely attractive and fundamentally different from other statistical arbitrage related concepts, the maximal loss is the initial investment $c\cdot S_{0}$ as the held strategy upon initialization, by its construction as a model-free hedge, never leads to negative payoffs. This means an investor can fix the level $c$ which imposes an upper bound on potential future losses. We illustrate the idea in Example 5.1 and Figure 7.
Example 5.1.
Consider a forward-start call option on Tesla with payoff
$\Phi(S_{t_{1}},S_{t_{2}})=(S_{t_{2}}-S_{t_{1}})^{+},$
where $S_{0}=800$ (Tesla’s spot price at initial time $t_{0}$), $t_{1}=$ March 1, and $t_{2}=$ March 29. Further, assume that the computed model-free bounds, computed via Algorithm 1, are given by
$\underline{\rm P}(\Phi)=34\quad\text{and}\quad\overline{\rm P}(\Phi)=44,$
and that the market price of the derivative is $\mathfrak{M}(\Phi)=42.5$, close to the upper bound. Then the trade is triggered since
$\min\{\mathfrak{M}(\Phi)-\underline{\rm P}(\Phi),\overline{\rm P}(\Phi)-\mathfrak{M}(\Phi)\}=1.5\leq c\cdot S_{0},$
for a threshold $c=0.02$. We thus invest $\overline{C}=\overline{\rm P}(\Phi)-\mathfrak{M}(\Phi)=1.5$ in the model-free super-hedge, which dominates the derivative payoff path-wise. Assume that on maturity we observe $S_{t_{1}}=810$ and $S_{t_{2}}=870$, so that:
$\Phi(S_{t_{1}},S_{t_{2}})=(870-810)^{+}=60.$
Note that, by construction, the model-free super-hedge leads a payoff that dominates the derivative payoff $\Phi(S_{t_{1}},S_{t_{2}})$ by a margin of at least $1.5$ with probability $\overline{q}(c)$. If we now assume that the model-free super-hedge realizes a payoff of $\overline{u}(S_{t_{1}},S_{t_{2}})=63$, then the payoff of our strategy is
$\overline{u}(S_{t_{1}},S_{t_{2}})-\Phi(S_{t_{1}},S_{t_{2}})=3,\quad\text{and net profit}=3-1.5=1.5.$
This leads in this example to a 100% return with a worst-case loss of at most $1.5$, highlighting the favorable risk–return profile of the strategy.
[…]
Figure 7. The figure illustrates the idea of the trading strategy: the strategy is initiated whenever the market price ${\mathfrak{M}}(\Phi)$ is close enough to one the price bounds $\overline{\mathrm{P}}(\Phi)$, $\underline{\mathrm{P}}(\Phi)$, illustrated via the blue shaded area. The probabilities $\underline{q}(\tfrac{{\mathfrak{M}}(\Phi)-\underline{\rm{P}}(\Phi)}{S_{0}})$ and $\overline{q}(\tfrac{\overline{\rm{P}}(\Phi)-{\mathfrak{M}}(\Phi)}{S_{0}})$ describe the probability for a profitable trade in dependence of the market price.
The choice of $c$ allows to steer the level of risk entailed in the strategy: namely, the larger $c$, the less restrictive is condition (5.1), and hence more trades are initiated coming however at the cost of a reduced probability of a profitable outcome (since $\underline{q}(\cdot)$ and $\overline{q}(\cdot)$ are decreasing functions). Moreover, the possible initial prices of the strategy increases with $c$, and therefore also the associated potential losses, implying an increased downside deviation of the outcomes. Finally, we remark that even though the estimation of the empirical measure $\widehat{\mathbb{P}}$ introduces model risk, the maximum downturn $c\cdot S_{0}$ is independent of $\widehat{\mathbb{P}}$ as it is based on a model-free trading strategy, i.e., since $\Phi(S)-\underline{u_{d,(\theta_{i,j}),(H_{i})}}(S)\geq 0$, and $\overline{u_{d,(\theta_{i,j}),(H_{i})}}(S)-\Phi(S)\geq 0$ pathwise, i.e., for all $S\in\mathbb{R}^{n}$.
5.2. Testing the strategy
To empirically test the strategy described in Section 5.1, we use the data outlined in Section 2 and split it into a training period (to construct $\widehat{\mathbb{P}}$), and a testing period (to backtest the strategy). All the derivatives considered for super- and sub- hedging in Section 4 are included for the empirical study of the strategy. The data related to the period from Jan $2018$ to Sep $2020$ is used as training period, the second half (Oct $2020$ to Dec $2022$) serves as test period.
Since we do not have access to historical prices of OTC-traded exotic derivatives, as a realistic surrogate we simulate derivative prices $\mathfrak{M}(\Phi)$ by a calibrated Heston-model as described in Appendix D. Due to the model-free construction, using derivative prices from a Heston model instead of market quotes only changes the number of trades being triggered but not the average profit conditional on trades being initiated.
Figure 8. Left: The graphs show $\underline{q}(c)$ and the the average returns of $\frac{\Phi(S)-\underline{u_{d,(\theta_{i,j}),(H_{i})}}(S)-\underline{C}}{S_{0}}$ conditional on $\frac{\mathfrak{M}(\Phi)-\underline{\operatorname{P}}}{S_{0}}\leq c$ in dependence of $c$, cost $\underline{C}=\underline{\operatorname{P}}-\mathfrak{M}(\Phi)$, evaluated on training data.
Right: The graphs show $\overline{q}(c)$ and the the average returns of $\frac{\overline{u_{d,(\theta_{i,j}),(H_{i})}}(S)-\Phi(S)-\overline{C}}{S_{0}}$ conditional on $\frac{\overline{\operatorname{P}}-\mathfrak{M}(\Phi)}{S_{0}}\leq c$ in dependence of $c$, cost $\overline{C}=\mathfrak{M}(\Phi)-\overline{\operatorname{P}}$, evaluated on training data.
Figure 8 illustrates the dependence of $\underline{q}(c)$ and $\overline{q}(c)$ on $c$ as well as the average returns of $\frac{\Phi(S)-\underline{u_{d,(\theta_{i,j}),(H_{i})}}(S)-\underline{C}}{S_{0}}$ conditional on ${\mathfrak{M}(\Phi)-\underline{\operatorname{P}}(\Phi)}\leq c\cdot S_{0}$, i.e., conditional on a trade being triggered as the market price is close enough to the lower model-free price bound, where we denote the costs of this strategy by $\underline{C}:=\mathfrak{M}(\Phi)-\underline{\operatorname{P}}(\Phi)$. Analogue, we study $\frac{\overline{u_{d,(\theta_{i,j}),(H_{i})}}(S)-\Phi(S)-\overline{C}}{S_{0}}$ conditional on ${\overline{\operatorname{P}}(\Phi)-\mathfrak{M}(\Phi)}\leq c\cdot S_{0}$ corresponding to the situation that a trade is triggered since the market price is close enough to the upper bound, where $\overline{C}:=\overline{\operatorname{P}}(\Phi)-\mathfrak{M}(\Phi)$. We evaluate on the training data. The two subplots at the bottom of Figure 8 show how the average net profit varies when gradually increasing $c$. Note that in this consideration the costs of the strategy are taken into account which explains the observation that with an increasing $c$ the average profit decreases due to the increasing costs of the strategy.
In Table 2 and Table 3 we illustrate the results of the suggested strategy evaluated on training data, where in contrast to Figure 8 we study annualized net profits per trade (no division by the spot value). By construction, both $\overline{q}(c)$ and $\underline{q}(c)$ decrease when $c$ increases and hence, combined with the increased costs of the strategy, the average return of the respective strategies decreases. Naturally, more strategies fulfil the selection criterion (5.1) with increasing $c$, implying that more trades are executed. An increasing $c$ also comes with a decreasing risk-adjusted return (in terms of Sortino ratio and Sharpe ratio). Also note that, since the strategies are almost not prone to losses (up to losing the initial investment), the variance is mainly driven by upward movements as being reflected in very high Sortino ratios.
Table 2. Train period performance results. Only trades triggered by $\overline{\operatorname{P}}(\Phi)-\mathfrak{M}(\Phi)\leq c\cdot S_{0},$ are considered.
$\overline{q}(c)$ $0.99$ $0.95$ $0.90$ $0.80$ $0.70$ $0.60$ $0.50$
c -0.0047 -0.0001 0.0014 0.0060 0.0123 0.0200 0.0300
Number of Trades 637 1142 1420 2653 4376 6502 9211
Avg return p.a. 2855.59 2350.08 2038.54 1293.50 858.34 632.62 447.80
Std dev. p.a. 367.10 593.99 536.26 408.17 326.35 327.66 276.96
Down dev. p.a. 7.43 31.14 15.88 20.47 24.66 26.36 26.90
Sharpe ratio 7.78 3.96 3.80 3.17 2.63 1.93 1.62
Sortino ratio 384.35 75.47 128.34 63.20 34.80 24.00 16.64
Max. Loss 21.11 22.58 22.58 24.51 52.14 67.00 143.69
Max. Profit 1073.39 3289.63 3289.63 3289.63 3289.63 13854.70 13854.70
We observe in Table 3 that even for relative large levels of $c$ only a small number of trades are executed because prices of the calibrated Heston model are much closer to the model-free upper bound, emphasizing the importance of including both lower bound and upper bound into the selection criterion (5.1).
Table 3. Train period performance results. Only trades triggered by $\mathfrak{M}(\Phi)-\underline{\operatorname{P}}(\Phi)\leq c\cdot S_{0}$ are considered.
$\underline{q}(c)$ $0.99$ $0.95$ $0.90$ $0.80$ $0.70$ $0.60$ $0.50$
c -0.0147 -0.0082 -0.0049 0.0009 0.0012 0.0049 0.0108
Number of Trades 0 0 0 2 2 4 11
Avg return p.a. -15.14 -15.14 -27.83 129.35
Std dev. p.a. 1.80 1.80 1.39 28.64
Down dev. p.a. 0.19 1.25
Sharpe ratio -8.42 -8.42 -20.02 4.52
Sortino ratio -146.13 103.86
Max. Loss 2.94 2.94 20.46 20.46
Max. Profit 9.97 9.97 9.97 374.60
Next, we evaluate the performance of the strategy for different choices of $c$ on testing data and illustrate the results in Table 4. The results confirm the profitability of the strategy, and they indicate that, as expected, in practice smaller levels of $c$ lead to higher risk-adjusted returns which however comes at the expense of trading less frequently, and therefore a restricted scalability of the strategy, which might in the extreme case even lead to no trades being triggered at all when choosing $c$ too small in an arbitrage-free market.
Table 4. Test period performance
$c$ 0% 1% 2.5% 5% 10%
$\overline{q}(c)$ 0.95 0.85 0.62 0.31 0.09
$\underline{q}(c)$ 1.00 0.99 0.94 0.15 0.14
Number of Trades 1346 4034 11848 41972 48211
Avg return p.a. 3059.80 133.84 667.53 327.86 336.49
Std dev. p.a. 577.92 382.87 310.50 304.07 530.71
Down dev. p.a. 4.78 35.49 51.42 77.43 72.23
Sharpe ratio 5.29 3.48 2.15 1.08 0.63
Sortino ratio 640.15 37.58 12.98 4.23 4.66
Max. Loss 28.44 62.68 111.82 187.33 305.37
Max. Profit 1675.88 1675.88 4120.82 15308.64 160994.61
6. Conclusion
In our study we have investigated the quality of model-independent super-hedging (and sub-hedging) strategies using data of liquidly traded options of constituents of the S&P 500 index. Our study reveals that model-free pricing and hedging approaches lead, compared with model-based hedges based on Black–Scholes, local volatility or Heston models, (see Section 4.3) to competitive hedging results even though, by construction as worst-case approaches they are known to be conservative.
Our quantification of the distribution of hedging gaps of model-free hedging strategies based on real data of the S&P 500 index across moneyness and different maturities shows that these model-free hedges often lead to payoffs closely aligned with actual realized payoffs of the underlying derivatives. Building on these findings, we present in Section 5 a novel highly profitable low risk trading strategy that exploits the empirical tightness of model-free bounds by trading in the model-free hedging whenever the market price is close enough to the theoretical price.
Our findings indicate that model-free methods, traditionally viewed as overly conservative, can in fact yield efficient and statistically robust trading strategies given the underlying option data is sufficiently liquid and hence informative enough. This finding opens door to broader practical adaption of robust approaches in the finance industry.
Appendix A Numerical Algorithm
In this section we provide with Algorithm 1 a detailed description of the numerical algorithm used to compute model-free price bounds.
Input : Number of traded call and put options $M_{1}$; bid and -ask prices of traded call options; interest rate $r$; Payoff function $\Phi$; Tolerance level $\operatorname{TOL}$; Large grid $\mathbb{G}\subset\mathbb{R}^{2}$ with number of meshes $N^{2}$; Initial sub-grid $\mathbb{G}_{0}\subset\mathbb{G}$;
for $k=0,1,\cdots$ do
   Find feasible parameters $x^{(k)}:=\left(d^{(k)},(\theta_{i,j}^{(k)})_{i=1,2,\atop j=1,\dots,M_{1}},H_{0}^{(k)},({H_{1}^{\ell}}^{(k)})_{\ell=1,\dots,N}\right)$ that minimize (3.5) on the grid $\mathbb{G}_{k}$, i.e.,
$\widetilde{\overline{\operatorname{P}}}(\Phi,\mathbb{G}_{k})=d^{(k)}+\sum_{i=1}^{2}\sum_{j=1}^{M_{i}}({\theta_{i,j}^{\rm call,+}}^{(k)}h_{i,j}^{\rm call,+}-{\theta_{i,j}^{\rm call,-}}^{(k)}h_{i,j}^{\rm call,-})$
with $u_{x^{(k)}}(S)\geq\Phi(S)\text{ for all }S\in\mathbb{G}_{k}$ for $u_{x^{(k)}}$ as defined in (3.1);
  Compute the constraint violation on $\mathbb{G}$ given by $\delta:=\min_{S\in\mathbb{G}}\{u_{x^{(k)}}(S)-\Phi(S)\}$;
 if $\delta>-\operatorname{TOL}$ then STOP;
 else
     Define $\mathbb{G}_{k+1}:=\mathbb{G}_{k}\cup\{S\in\mathbb{G}\leavevmode\nobreak\ |\leavevmode\nobreak\ u_{x^{(k)}}(S)-\Phi(S)<\delta+\frac{1}{k+1}\};$
   end if
end for
Output : Parameters $d^{(k)},(\theta_{i,j}^{(k)})_{i=1,2,\atop j=1,\dots,M_{1}},H_{0}^{(k)},({H_{1}^{\ell}}^{(k)})_{\ell=1,\dots,N}$;
Algorithm 1 Cutting plane algorithm to compute super-hedging strategies
Appendix B Hyperparameter - Selection
B.1. Parameter optimization
Applying Algorithm 1 requires selecting hyperparameters that are used as input to the algorithm. To find best-working hyperparameters and understand its influence on the outcomes, we conducted a systematic exploration of the involved hyperparameters, involving the parameters tolerance ($\rm TOL$), initial grid size ($n$), and large grid size ($N$).
B.1.1. Superhedging gap and cutting-plane algorithm
For our analysis, Algorithm 1 was applied to compute the gap between the superhedging value and the actual hedge performance on a subset of the dataset described in Section 2.
To understand the effects of the parameters on the algorithm’s performance, we visualize in Figures 9 and 10 the effects of varying the parameters on the total time taken for the cutting plane algorithm to converge (Runtime), the number of iterations required for convergence (Iteration count) and on the relative size of the difference between the superhedging value and actual hedge (Gap percentage).
Figure 9. Left: The plot depicts how different grid sizes influence the computational time, Right: The plot illustrates the interaction between grid sizes (NN) and the number of iterations ($n$) required for convergence (reaching the tolerance level).
Figure 10. The plots illustrate the impact of grid sizes (NN) and the number of iterations ($n$) on the gap between superhedging and actual payoff of the derivative.
These visualizations helped us identify optimal parameter combinations that minimize both the gap and runtime while ensuring sufficient convergence. Details of the used hyperparameters can be found under https://github.com/QIYIHAN/Empirical-Price-Bounds.
Appendix C Distribution of the payoff gaps
Here, in Figure 1114 we provide histograms illustrating the distributions of the payoff gaps of the experiments from Section 4.1.
Figure 11. The histogram of the hedging gap with payoff function $(S_{t_{2}}-S_{t_{1}})^{+}$.
Figure 12. The histogramof the hedging gap with payoff function $\frac{S_{t_{1}}+S_{t_{2}}}{2}-\sqrt{S_{t_{1}}\cdot S_{t_{2}}}$.
Figure 13. The histogram of the hedging gap with payoff function $\log(S_{t_{2}}/S_{t_{1}})$.
Figure 14. The histogram of the hedging gap with payoff function $\sqrt{S_{t_{1}}\cdot S_{t_{2}}}$.
Appendix D Benchmark Models
In this section we briefly describe the benchmark models used in Section 4.3 and the method to compute prices of forward start option and, based on this, its delta sensitivities. For all of the models, we approximate the option delta as $\Delta(S_{t},t):=\frac{{\rm P}\left(S_{t}\cdot(1+\alpha),t\right))-{\rm P}(S_{t},t)}{\alpha S_{t}}$ on a daily basis, where we use $\alpha=0.001$, and where ${\rm P}(S_{t},t)$ denotes the option price in the respective model at time $t$ when the underlying security attains value $S_{t}$.
Black–Scholes model
The Black–Scholes model (black1973pricing) models the evolution of the asset value $S_{t}$ by the stochastic differential equation
${\mathop{}\!\mathrm{d}}S_{t}$ $=rS_{t}{\mathop{}\!\mathrm{d}}t+\sigma S_{t}{\mathop{}\!\mathrm{d}}W_{t}$
for some $r\in\mathbb{R}$, $\sigma>0$ and some standard Brownian motion $(W_{t})_{t\geq 0}$. We use the same interest rate $r$ as in the model-free considerations, described in Section 3.3.2 as well as the Black–Scholes implied volatility $\sigma^{\rm BS}$, i.e., the volatility $\sigma^{\rm BS}$ that makes the call option market prices match the model-based call option prices. For the Black–Scholes model, the option price $P(S,t)$ is calculated as the analytical price of the forward-start option price formula with $k=1$ in (musiela1997martingale, p. 231).
Heston model
As a second benchmark model we consider the stochastic volatility model introduced by heston1993closed where the stock price $(S_{t})_{t\geq 0}$ in the continuous-time Heston model is given by the stochastic differential equations
${\mathop{}\!\mathrm{d}}S_{t}$ $=rS_{t}{\mathop{}\!\mathrm{d}}t+\sqrt{\nu_{t}}S_{t}{\mathop{}\!\mathrm{d}}W_{t}^{1},$
${\mathop{}\!\mathrm{d}}\nu_{t}$ $=\kappa(\theta-\nu_{t}){\mathop{}\!\mathrm{d}}t+\sigma\sqrt{v_{t}}{\mathop{}\!\mathrm{d}}W_{t}^{2},$
for $r$ being the interest rate, $\kappa$ the volatility mean-reversion speed, $\theta$ the volatility long term mean, $\sigma$ the diffusion parameter associated to the volatility, and where $W_{t}^{1}$ and $W_{t}^{2}$ are two correlated Brownian motions. The parameters of the Heston-model are calibrated to the considered call option prices according to the methodology described in levendorskiui2012efficient. The price of the forward start option (and based on this its sensitivities) can be computed as described in Kruse2005.
Dupire’s Local Volatility model
Dupire’s local volatility model (dupire1994pricing and derman1994riding) model the asset price $(S_{t})_{t\geq 0}$ via
${\mathop{}\!\mathrm{d}}S_{t}$ $=rS_{t}{\mathop{}\!\mathrm{d}}t+\sigma(t,S_{t})S_{t}{\mathop{}\!\mathrm{d}}W_{t}$
for interest rate $r$ and Brownian motion $(W_{t})_{t\geq 0}$ as well as
(D.1) $\sigma^{2}(T,K)=2\frac{\frac{\partial}{\partial T}c(T,K)+Kr\frac{\partial}{\partial K}c(T,K)}{K^{2}\frac{\partial^{2}}{\partial K^{2}}c(T,K)}$
with $K$ and $T$ being strike and maturity, respectively of the considered call option with price $c(T,K)$. The derivatives in the right-hand-side of (D.1) are approximated via finite differences.
The delta sensitivity is then derived from a corresponding Black–Scholes model, considering the option price as a function of stock price $S$ with volatility $\sigma$, $P=P(S,\sigma(t,S))$. The option delta can then be calculated analytically as
$\frac{dP}{dS}=\frac{\partial P}{\partial S}\bigg{|}_{\sigma}+\frac{\partial P}{\partial\sigma}\bigg{|}_{S}\frac{\partial\sigma(T,K,S)}{\partial S}=\Delta_{\text{BS}}(S,\sigma)+\nu_{\text{BS}}(S,\sigma)\frac{\partial\sigma(T,K)}{\partial S}$
for $\nu_{\text{BS}}$ and $\nu_{\text{BS}}$ denoting Delta and Vega, respectively in the Black–Scholes model.